Solve to find x. 2x+y=3 and -7x+2y=17.
Answer by lawdej(58)

2x+y=3
-7x+2y=17.
to solve this multiply 2x+y=3 by -2 to cancel y
so here is the result
-4x-2y=-6
-7x+2y=17
now add both of them the result is
-11x=11
x=-11/11
x=-1 so this is the value of x.
but if you need to find the value of y substitute x=-1 to either of 2x+y=3 or -7x+2y=17
for example take 2x+y=3
2(-1)+y=3
-2+y=3
y=3+2=5